# A Trifecta of Modelling Tools: Navigating COMPAS, Shark and Bayesian Inference for Binary Black Hole Model Selection

### v1

@@ 5.2 Bayes Factor and Likelihood Analysis
We note that the Bayes factor indicates there is strong evidence that some groups of models are preferred more than others, which we will further discuss in the likelihood analysis. We emphasise these with highlighted boxes in Fig. 12 labeled A-G. We can understand the range of Bayes factor by considering the different components that enter Eq 33. The difference in $\log\beta(\lambda)$ between different models ranges from $\sim 0.05-3.44$. However, we find that the $I^{k}$ terms varies over 50 orders of magnitude (see bottom plot in Fig. 18) and therefore dominates the Bayes factor. This confirms that some of the models are distinguishable and strongly preferred (more likely to produce progenitors of particular GW events) over others purely based on the masses and redshift distributions they predict. When including $N_{\lambda}$ as a parameter in the likelihood, the $-\mu_{\lambda}+N_{\rm obs}\log(\mu_{\lambda})$ terms this introduces into the log-likelihood has significant weight, and comes to dominate the Bayes factor, meaning that a comparison of two models using this likelihood strongly favours those that produce a number of events close to the observed number of events in our data sample, even if the event-level distributions are less well matched. Hence, we carefully consider the results using both of our likelihoods.
@@
Having established groups of models that are strongly preferred over other models, we now discuss the likelihoods for individual cases and the trends within these groups. The following results discussed use the $\rm SNR\geq 8$ selection cut. We see an increase in the likelihood across all models, but the trends don’t change substantially with the selection criteria and we find the differences are made more distinct when using the $\rm SNR\geq 8$ sample. We show this in Tables 4, 5 and 6 with Bayes factors between our various models.
@@
We see an increase in the likelihood from $f_{\gamma}=0$ to $f_{\gamma}=0.5$ and we find that the likelihoods are similar with no CHE and pessimistic and optimistic CHE. In Fig. 13, we show the likelihoods for fixed $f_{\gamma}$. It is clear that $f_{\gamma}=0.5$ is preferred over $f_{\gamma}=1.0$. Less BBH mergers are produced with $f_{\gamma}=1.0$. Models with $f_{\gamma}=0.5$ produce more equal mass binaries, which is more consistent with the LVK population (Abbott et al., 2019b). However, $f_{\gamma}=0.75$ is unique in that for low mass transfer $f_{\rm MT}<0.5$ the population is disfavoured similar to $f_{\gamma}=1.0$, but for high mass transfer $f_{\rm MT}>0.5$ the model is more favourable. In all cases, we see little impact from the choice of common envelope efficiency. This highlights the degeneracy between stellar modelling assumptions and populations. Ultimately, we find that in order to reproduce the observed events from LVK, we need either momentum loss to occur close to the accretor, or, if occurring further away, for the mass transfer fraction to be high. These trends are also highlighted by boxes A and B in Fig. 12, where moderate values of $f_{\gamma}$ are strongly favoured when comparing population parameters.
@@
Figure 13: Likelihoods for models with no CHE. Each band corresponds to a fixed $f_{\gamma}$. Within each band, excluding the Rauf et al. (2023) models in blue, the mass transfer fraction and common envelope efficiency are increasing with increasing model number, within the ranges in Table 1, with $f_{\rm MT}$ varying more rapidly.
@@ 5.2.1 Likelihood including $N_{\lambda}$
Figure 15: Model versus unmarginalised likelihood, where each coloured band corresponds to the a fixed CHE mode and $f_{\gamma}$. Within each band, excluding the Rauf et al. (2023) models, the mass transfer fraction and common envelope efficiency are increasing with increasing model number, within the ranges in Table 1, with $f_{\rm MT}$ varying more rapidly.
@@ 5.3 Which data are the most informative?
This is particularly emphasised for GW190521 where we found that the posterior samples often did not fall at all within the chirp mass or mass ratio range output from COMPAS. In these cases we assign a probability of zero to these events such that they make no difference to the sum in Eq 23 and 28, but are also hence not represented in Fig. 18. Model choices that fail to produce any progenitor close to GW190521 include:
@@ 6 Conclusion
The COMPAS population synthesis code contains a number of free parameters — we introduce some of the most salient of these in Section 2. Our results fitting these models to data in Section 5 demonstrate there is a preference for subsets of models dictated by the type of isolated formation channel and specific angular momentum loss. However, we frequently find opposing trends in the likelihoods with and without $N_{\lambda}$ dependence — models that produce realistic event-level masses and redshifts frequently over-predict the number of events we should have observed to date. Looking at models which perform well with both likelihoods, we find that models with no Chemically Homogeneous Evolution, low to moderate values of $f_{\gamma}$ (such that angular momentum loss occurs closer to the donor) and higher rates of mass-loss from Wolf–Rayet winds (larger $f_{\rm WR}$) are generally preferred by current data.
@@
We find that the selection effects impact our likelihoods, as the $\rm SNR\geq 8$ cut on the GW events increases the likelihood for all models. This may vary with future GW detections, particularly if the BH masses are in the mass gaps. We also use the individual event likelihoods $I^{k}$ across all our models term to distinguish individual events that are easy or more difficult to reproduce by our models. We identify the usual suspects GW190521 and GW190814 as typically having low probability of existing in our new models, as well as some outlier events in the the lower and upper mass gap. This motivates future studies to consider joint population inference and model selection studies across multiple formation channels (Gompertz et al., 2022; Gayathri et al., 2023).
@@
Our method for efficiently calculating the merger rate can also be applied to other population synthesis codes. However, we note that it is only applicable in the case where the metallicity dependent SFR is the same for the fiducial and new model. We argue that using a different SFR may not impact the PDF significantly. van Son et al. (2023) finds that variation in the SFR do not impact the features in the primary BH mass distribution and have minimal impact on the low mass end of the distribution.
@@
We acknowledge this work relies on various assumptions and simplifications. For future work, we consider including the effective spin as a population parameter in the analysis. The covariance between the mass ratio and effective spin can arise due to common envelope physics, mass transfer efficiency and Eddington limited accretion (Mandel & Fragos, 2020; Bavera et al., 2021; Zevin & Bavera, 2022; Broekgaarden et al., 2022b; Adamcewicz et al., 2023; Olejak et al., 2024). This could be used to further distinguish between COMPAS models. Alternatively, we can build the PDF only with $\mathcal{M}_{c}$, mass ratio or redshift to test the validity of the model outputs. We could also consider including natal kick velocity as one of the hyper-parameters to vary in COMPAS in the future, as this is not well constrained for BHs (Callister et al., 2021; Cui & Li, 2023; Vigna-Gómez et al., 2024), or more fundamental parameters such as the binary fraction. Varying the selection function cuts applied to current data to restrict the analysis to only the strongest detections may further inform us on trends but given we found consistent results with our two different selection functions this may not necessarily provide improved constraints on the hyper-parameters.
@@ Acknowledgements
We thank Carole Perigois and Michella Mapelli for useful discussion on their paper that motivated this work. We thank Ilya Mandel, Eric Thrane and Paul Lasky for useful discussions on the methodology and results. We also thank Reinhold Wilcox and Floor Broekgaarden for useful discussions on the COMPAS hyper-parameters and relevant literature.
@@ Appendix A Model initial conditions and fixed parameters
Distribution Model
IMF Kroupa (2002)
Initial mass range $5-150M_{\odot}$
$5-200M_{\odot}$ (Rauf et al., 2023)
Semi-major axis Flat-in-log
Orbital period Flat-in-log
Mass ratio Sana et al. (2012)
Eccentricity ZERO (circular)
Sana et al. (2012) (Rauf et al., 2023)
Metallicity Shark (Lagos et al., 2018)
Table 2: Summary of COMPAS default initial conditions.

### v2

@@ 5.2 Bayes Factor and Likelihood Analysis
We note that the Bayes factor indicates there is strong evidence that some groups of models are preferred more than others, which we will further discuss in the likelihood analysis. We emphasise these with highlighted boxes in Fig. 12 labeled A-G. We can understand the range of Bayes factor by considering the different components that enter Eq. 33. The difference in $\log\beta(\lambda)$ between different models ranges from $\sim 0.05-3.44$. However, we find that the $\mathcal{I}^{k}$ terms are distributed over 50 orders of magnitude (see bottom plot in Fig. 18) and therefore dominates the Bayes factor. This confirms that some of the models are distinguishable and strongly preferred (more likely to produce progenitors of particular GW events) over others purely based on the masses and redshift distributions they predict. When including $N_{\lambda}$ as a parameter in the likelihood, the $-\mu_{\lambda}+N_{\rm obs}\log(\mu_{\lambda})$ term this introduces into the log-likelihood has significant weight, and comes to dominate the Bayes factor, meaning that a comparison of two models using this likelihood strongly favours those that produce a number of events close to the observed number of events in our data sample, even if the event-level distributions are less well matched. Hence, we carefully consider the results using both of our likelihoods.
@@
Having established groups of models that are strongly preferred over other models, we now discuss the likelihoods for individual cases and the trends within these groups. The following results (Fig. 13, 14, 15, 16, 17, 18) discussed use the $\rm SNR\geq 8$ selection cut. We see an increase in the likelihood, due to an increase in $N_{\rm obs}$, across all models compared to using $p_{\rm astro}\geq 0.9$ and $\rm FAR\leq 0.25yr^{-1}$, but the trends don’t change substantially with the selection criteria and we find the differences are made more distinct when using the $\rm SNR\geq 8$ sample. We show this in Tables 4, 5 and 6 with Bayes factors between our various models.
@@
We see an increase in the likelihood from $f_{\gamma}=0$ to $f_{\gamma}=0.5$ and we find that the likelihoods are similar with no CHE and pessimistic and optimistic CHE. In Fig. 13, we show the likelihoods for fixed $f_{\gamma}$. It is clear that $f_{\gamma}=0.5$ is preferred over $f_{\gamma}=1.0$. Less BBH mergers are produced with $f_{\gamma}=1.0$. Models with $f_{\gamma}=0.5$ produce more equal mass binaries, which is more consistent with the LVK population (Abbott et al., 2019b). However, $f_{\gamma}=0.75$ is unique in that for low mass transfer $f_{\rm MT}<0.5$ the population is disfavoured similar to $f_{\gamma}=1.0$, but for high mass transfer $f_{\rm MT}>0.5$ the model is more favourable. In all cases, we see little impact from the choice of CE efficiency. This highlights the degeneracy between stellar modelling assumptions and populations. Ultimately, we find that in order to reproduce the observed events from LVK, we need either momentum loss to occur close to the accretor, or, if occurring further away, for the mass transfer fraction to be high. These trends are also highlighted by boxes A and B in Fig. 12, where moderate values of $f_{\gamma}$ are strongly favoured when comparing population parameters.
@@
Figure 13: Likelihoods for models with no CHE. Each band corresponds to a fixed $f_{\gamma}$. Within each band, excluding the Rauf et al. (2023) models in blue, the mass transfer fraction and CE efficiency are increasing with increasing model number, within the ranges in Table 1, with $f_{\rm MT}$ varying more rapidly.
@@ 5.2.1 Likelihood including $N_{\lambda}$
Figure 15: Model versus unmarginalised likelihood, where each coloured band corresponds to the a fixed CHE mode and $f_{\gamma}$. Within each band, excluding the Rauf et al. (2023) models, the mass transfer fraction and CE efficiency are increasing with increasing model number, within the ranges in Table 1, with $f_{\rm MT}$ varying more rapidly.
@@ 5.3 Which data are the most informative?
This is particularly emphasised for GW190521 where we find that the posterior samples often did not fall at all within the chirp mass or mass ratio range output from COMPAS. In these cases we assign a probability of zero to these events such that they make no difference to the sum in Eq. 23 and 28, but are also hence not represented in Fig. 18. Model choices that fail to produce any progenitor close to GW190521 include:
@@ 6 Conclusion
The COMPAS population synthesis code contains a number of free parameters — we introduce some of the most salient of these in Section 2. These are the mass transfer fraction, CE efficiency, specific AM loss fraction, CHE, remnant mass prescription and Wolf–Rayet winds . We find that various combinations of these hyper-parameters impact the chirp mass, mass ratio and redshift distribution, as well as changing the orbital separations of binaries, which affects the number of BBH mergers within an observed time. Our results fitting these models to data in Section 5 demonstrate there is a preference for subsets of models dictated by the type of isolated formation channel and specific AM loss. However, we frequently find opposing trends in the likelihoods with and without $N_{\lambda}$ dependence — models that produce realistic event-level masses and redshifts frequently over-predict the number of events we should have observed to date. Looking at models which perform well with both likelihoods, we find that models with no CHE, low to moderate values of $f_{\gamma}$ (such that AM loss occurs closer to the donor) and higher rates of mass-loss from Wolf–Rayet winds ($f_{\rm WR}>1$) are generally preferred by current data. We provide a quantitative summary of the hyper-parameters for the models that best reproduce the GW population excluding GW190521 at the end of Section 5.3.
@@
We find that the selection effects impact our likelihoods. Although the selection allows for potentially non-astrophysical sources in the sample, the $\rm SNR\geq 8$ cut on the GW events increases the likelihood for all models compared to $\rm FAR\leq 0.25yr^{-1}$ and $p_{\rm astro}\geq 0.9$. Preference for models do not change with either selection criteria. However, these conclusions may vary with future GW detections, particularly if the BH masses are in the mass gaps. We also use the individual event likelihoods $\mathcal{I}^{k}$ across all our models term to distinguish individual events that are easy or more difficult to reproduce by our models. We identify the usual suspects GW190521 and GW190814 as typically having low probability of existing in our new models, as well as some outlier events in the the lower and upper mass gap. This motivates future studies to consider joint population inference and model selection studies across multiple formation channels (Gompertz et al., 2022; Gayathri et al., 2023).
@@
Our method for efficiently calculating the merger rate can also be applied to other population synthesis codes. However, we note that it is only applicable in the case where the metallicity dependent SFR is the same for the fiducial and new model. We argue that using a different SFR may not impact the PDF significantly. van Son et al. (2023) finds that variation in the SFR do not impact the features in the primary BH mass distribution and have minimal impact on the low mass end of the distribution. For the redshift range studied in this paper, Shark predicts a metallicity dependent SFR density compared to the fiducial model explored by Neijssel et al. (2019), which produces a BBH merger rate comparable to GWTC-1. This may be contributing to our models overestimating the number of events. Variations in stellar evolution in population synthesis (Broekgaarden et al., 2022a; Chruślińska, 2024) and modelling of galaxy distributions in galaxy simulations [^11] We refer to Fig. 2 and 3 of Rauf et al. (2023) for comparisons of Shark to observations. (Furlong et al., 2015; Lagos et al., 2024) can translate to large uncertainties in the SFR, that will impact the BBH merger rate. These uncertainties are non-negligible and should be accounted for in future studies of BBH model selection.
@@
We acknowledge this work relies on various assumptions and simplifications. For future work, we consider including the effective spin as a population parameter in the analysis. The covariance between the mass ratio and effective spin can arise due to CE physics, mass transfer efficiency and Eddington limited accretion (Mandel & Fragos, 2020; Bavera et al., 2021; Zevin & Bavera, 2022; Broekgaarden et al., 2022b; Adamcewicz et al., 2023; Olejak et al., 2024). This could be used to further distinguish between COMPAS models. Alternatively, we can build the PDF only with $\mathcal{M}_{c}$, mass ratio or redshift to test the validity of the model outputs. We could also consider including natal kick velocity as one of the hyper-parameters to vary in COMPAS in the future, as this is not well constrained for BHs (Callister et al., 2021; Cui & Li, 2023; Vigna-Gómez et al., 2024), or more fundamental parameters such as the binary fraction. Varying the selection function cuts applied to current data to restrict the analysis to only the strongest detections may further inform us on trends but given we find consistent results with our two different selection functions this may not necessarily provide improved constraints on the hyper-parameters.
@@ Acknowledgements
We thank Carole Perigois and Michella Mapelli for useful discussion on their paper that motivated this work. We thank Ilya Mandel, Eric Thrane and Paul Lasky for useful discussions on the methodology and results. We also thank Floor Broekgaarden for useful discussions on the COMPAS hyper-parameters and relevant literature. We also thank the anonymous reviewer for their helpful comments, that improved the quality of the manuscript.
@@ Appendix A Model initial conditions and fixed parameters
Distribution Model Range
IMF Kroupa (2002) $[5,150]M_{\odot}$
$[5,200]M_{\odot}$ (Rauf et al., 2023)
Semi-major axis Flat-in-log $[0.01,1000]$ AU
Orbital period Flat-in-log $[0.1,1000]$ days
Mass ratio Sana et al. (2012) $[0.01,1]$
Eccentricity ZERO (circular) 0
Sana et al. (2012) (Rauf et al., 2023) $[0,1]$
Log Metallicity Shark (Lagos et al., 2018) $[-7.0,-0.3]$
Table 2: Summary of COMPAS default initial conditions and range.